OpenStudy (anonymous):

PLEASE HELP! Consider c = 5200 and d = 0.0000037 Write down the value of r = c × d . Write down your value of r in the form a ×10k , where 1≤ a <10 and k ∈ Z

OpenStudy (anonymous):

\[5.2\times 10^3\times 3.7\times 10^{-6}\]

OpenStudy (anonymous):

\[=0.01924=1.924\times 10^{-2}\]